# Dynamic Benchmark Construction for Evaluating Large Language Models on Real-World Codes

### v1

@@ Abstract
As large language models (LLMs) become increasingly integrated into software development workflows, rigorously evaluating their performance on complex, real-world code generation tasks has become essential. However, existing benchmarks often suffer from data contamination and limited test rigor, constraining their ability to reveal model failures effectively. To address these, we present Code2Bench, a systematic, end-to-end pipeline for dynamically constructing robust and contamination-resistant benchmarks from real-world GitHub repositories. Specifically, Code2Bench introduces three key innovations: (1) Automated Dynamism, achieved through periodic ingestion of recent code to minimize training data contamination; (2) Scope Graph-based dependency analysis, which enables structured classification of functions into benchmark instances with controlled dependency levels (distinguishing between Self-Contained (SC) tasks for cross-language evaluation and Weakly Self-Contained (WSC) tasks involving permitted library usage); and (3) Property-Based Testing (PBT) for the automated synthesis of rigorous, high-coverage test suites to enable thorough functional verification. Using this pipeline, we construct Code2Bench-2505, the first dynamic benchmark derived from 880 recent Python projects spanning diverse domains, comprising 1,163 code generation tasks with 100% average branch coverage on ground-truth implementations. Extensive evaluation of 16 LLMs using Code2Bench-2505 reveals that models consistently struggle with SC tasks requiring complex, non-standard logic and cross-language transfer, while showing relatively stronger performance on WSC tasks in Python. Our work introduces a contamination-resistant, language-agnostic methodology for dynamic benchmark construction, offering a principled foundation for the comprehensive and realistic evaluation of LLMs on real-world software development tasks. The codes, data and results can be found at https://code2bench.github.io/.
@@ 1 Introduction
Recently, Large Language Models (LLMs) have been increasingly integrated into software development workflows [1, 2, 3], such as Github Copilot[4] and Cursor[5]. Therefore, the accurate and realistic evaluation of their performance on complex and real-world code generation tasks is essential [6, 7, 8, 9]. This necessitates robust benchmarks capable of capturing the complexities of real-world coding tasks while providing reliable and scalable evaluation.
However, existing methods individually address only subsets of these issues, and suffer from data contamination and limited test rigor. In particular, the primary challenge ❶ is the potential for training data contamination [10, 11, 12, 13], which compromises evaluation validity. Most benchmarks  [14, 15, 16] rely on static datasets, and their static nature makes them highly susceptible to contamination as LLMs are trained on large code corpora, including publicly available benchmark data. Although dynamic “live” benchmarks [17, 18, 19, 20] have emerged (e.g., using competitive programming or user interactions), they often do not extract tasks directly from the full complexity of real-world software repositories. Meanwhile, the challenge ❷, as highlighted by EvalPlus[21], a pervasive issue with many benchmarks is their insufficient test rigor, which fundamentally constrains their ability to effectively reveal subtle model failures. Most benchmarks employ a limited number of cases that are typically manually-crafted or generated using less rigorous methods [22, 23]. Even benchmarks derived from real-world code or repository contexts  [24, 25, 20, 26, 27], often lack comprehensive, automated, rigorous testing pipelines, relying instead on limited execution-based tests or manual verification. As summarized in Table 1, existing methods fall short across key dimensions, highlighting the significant limitations remaining and the necessity to design new benchmarks.
@@
Table 1: Comparison of Code2Bench-2505 with existing code generation benchmarks.
@@
HumanEval [14] Manual
MBPP [15] Manual
EvalPlus [21] Manual
LiveCodeBench [19] Contests
RepoBench [24] Project Codebases
HumanEval-X [28] Manual
BigCodeBench [23] Synthetic
DevEval [26] Project Codebases
EvoCodeBench [20] Project Codebases
Code2Bench-2505 (Ours) Project Codebases
To address these challenges, we present Code2Bench, a novel, automated, end-to-end framework for dynamically constructing rigorous, contamination-resistant benchmark instances directly from recent real-world GitHub repositories. Code2Bench uniquely integrates key properties missing in existing approaches, offering: (1) Automated Dynamism, achieved through continuously ingest code from active GitHub repositories, focusing on recent commits. This periodic update mechanism is fundamental to minimizing training data contamination risk inherent in static benchmarks and ensures the benchmark remains relevant to the evolving software development ecosystem. (2) Scope Graph-based dependency analysis, which enables structured classification of functions into instances with controlled dependency levels (distinguishing between Self-Contained (SC) tasks for cross-language evaluation and Weakly Self-Contained (WSC) tasks involving controlled usage of specific, commonly used external libraries); and (3) Property-Based Testing(PBT) for the automated synthesis of rigorous, high-coverage test suites by dynamically generating diverse inputs, which enables the discovery of subtle errors and edge cases that limited test cases might miss. Furthermore, the language-agnostic Scope Graph analysis and the broad availability of PBT libraries across common programming languages collectively enable the multi-language extensibility. of Code2Bench.
Using Code2Bench, we construct Code2Bench-2505, the first dynamic benchmark instance derived from recent real-world Python projects. Code2Bench-2505 features 1163 code generation tasks across diverse domains, with controlled dependencies (SC/WSC) and rigorous PBT-generated tests achieving 100% branch coverage on ground-truth implementations. Evaluation of 16 state-of-the-art LLMs using Code2Bench-2505 reveals that models consistently struggle with SC tasks requiring complex and non-standard logic, while showing relatively stronger performance on WSC tasks in Python. Furthermore, we investigate the nature of LLM failures, demonstrate the effectiveness of rigorous PBT tests and real-world code source in revealing nuanced model limitations that might be missed by less challenging or static benchmarks. We summarize our contributions as follows:
  • We propose Code2Bench, a novel end-to-end pipeline for dynamically constructing rigorous benchmarks from recent real-world code, directly addressing contamination, realism, dependency control, test rigor, and construction scalability.
  • Using Code2Bench, we construct the first benchmark instance Code2Bench-2505, a high-quality dataset derived from authentic code with controlled dependencies and rigorous PBT tests.
  • We provide key insights into state-of-the-art LLM performance on Code2Bench-2505, revealing cross-language inconsistencies and surprising performance trend between SC and WSC tasks.
2 The Code2Bench Framework
@@
The Code2Bench framework is an automated and end-to-end pipeline meticulously designed to address the multifaceted challenges in evaluating LLMs on complex, real-world code tasks, including data contamination, insufficient rigor, dependency handling, and multi-language support. As illustrated in Figure 1, Code2Bench operates in two main stages: Candidate Filtering and Benchmark Instance Construction, which transform raw code into structured benchmark instances ready for evaluation. The design of this pipeline is underpinned by three foundational pillars:
  • Automated Dynamism: To mitigate data contamination risks, the framework continuously ingests code from active GitHub repositories, focusing on recent updates. This automated process ensures the benchmark evolves with the real-world software ecosystem, providing fresh evaluation tasks.
  • Rigorous Control: Achieving reliable evaluation requires rigorous testing and controlled task complexity. This pillar is implemented across two stages:
    • Within the Function Filtering stage, Scope Graph-based dependency analysis systematically identify and classify functions. This process enables the creation of task instances with controlled dependency levels (SC and WSC), ensuring realistic yet manageable complexity.
    • The Benchmark Instance Construction stage employs Property-Based Testing (PBT) for automated test suite generation. PBT’s ability to synthesize diverse, constraint-aware inputs based on function properties ensures rigorous, high-coverage test cases that effectively reveal subtle functional errors and edge cases. An LLM is considered to have successfully solved a problem instance if it passes all PBT-generated tests.
  • Language-Agnostic Core: The framework features a language-agnostic core, designed to ensure comprehensive multi-language extensibility. Key components like the Scope Graph analysis, based on universal program concepts, and the Property-Based Testing methodology, applicable across languages with PBT libraries, exemplify this design principle.
Crucially, the end-to-end pipeline rigorously filters and subjects function candidates to strict quality controls in test generation, ensuring that only high-quality, challenging, and well-tested instances derived from real-world code are included in the final benchmark. Given the immense scale and continuous evolution of platforms like GitHub, this rigorous process is capable of identifying and extracting a sufficient pool of such instances for benchmark construction at scale.
2.1 Candidate Filtering
The Function Filtering stage is designed to systematically extract suitable function candidates from the vast and often noisy landscape of real-world code repositories (Figure 1, left). This process is crucial for addressing challenges like data contamination and ensuring task realism and testability. It involves a multi-step pipeline employing various analysis techniques to select functions that are realistic, non-trivial, testable, and possess controllable dependencies.
Function Recall and Temporal Filtering
To obtain authentic tasks and mitigate contamination risk, we extract functions directly from actively developed GitHub codebases. We prioritize recent code by monitoring designated repositories and analyzing recent commit history. This temporal filtering helps exclude code potentially present in models’ pre-training data.
Preprocessing and Data Structuring
Following function identification, the relevant source code is preprocessed using Tree-sitter [29] to build Abstract Syntax Trees (ASTs). From these ASTs, essential relational information for each function, such as source code, parameters, type hints, and import statements, is extracted. This step standardizes the code representation and facilitates efficient subsequent analysis and filtering based on code structure and metadata.
Scope Graph based Dependency Analysis
A key requirement for benchmark tasks is having controllable dependencies to ensure testability and reduce contamination vectors from complex project structures. Using a language-agnostic Scope Graph based reference resolution technique [30], we analyze external dependencies for each function. This method identifies “unresolved references”—identifiers used but not defined locally. Based on this analysis, functions are classified: Self-Contained (SC) functions have no unresolved external references beyond language primitives, while Weakly Self-Contained (WSC) functions rely only on a predefined set of allowed, common libraries ($L_{allowed}$). Functions with dependencies outside $L_{allowed}$ (e.g., custom project modules) are discarded, ensuring selected candidates are either independent or have manageable, common dependencies crucial for reliable and standardized evaluation. Details can be found in Appendix B.
Program Analysis for Testability and Complexity
To ensure candidate functions are functionally testable and non-trivial, we apply program analysis, details of which can be found in Appendix C. We use Control Flow Graph (CFG) analysis to filter functions lacking verifiable output (e.g., no returns or only constants). Task complexity is measured via Cyclomatic Complexity [31]. Functions whose complexity does not fall within a predefined range are excluded from the analysis. To avoid duplicates, we perform function deduplication using exact function names and AST matching.
LLM-based Semantic Filtering
Structural and dependency analysis ensures technical feasibility, but may include semantically trivial functions like simple getters/setters. To ensure the final benchmark consists of challenging, relevant programming tasks, we employ an LLM-as-a-judge [32, 33] in a final filtering stage. A capable LLM assesses the semantic relevance and complexity of remaining candidates. Trivial functions identified by the LLM judge are discarded.
2.2 Benchmark Instance Construction
This stage transforms filtered function candidates into standardized benchmark instances, each accompanied by a comprehensive test case set, an executable test runner, and a precise instruction.
Test Case Generation via Property-Based Testing
A cornerstone of Code2Bench is the generation of high-quality, rigorous test cases using PBT (Figure 1, middle PBT detail). Unlike traditional example-based testing, which relies on fixed, pre-defined input-output pairs, PBT defines abstract properties that a piece of code should satisfy for a wide range of valid inputs. For code generation evaluation, the core property we test is functional equivalence: for any valid input, the output of the generated function must match the output of the ground-truth function implementation. We leverage PBT libraries (e.g., Hypothesis for Python) that allow us to define strategies for generating diverse inputs based on the function’s parameter types and constraints inferred from the ground-truth code or its specification (e.g., type hints, refined docstrings). The strategy builder process systematically explores the input space guided by these strategies and constraints. For each input $x_{i}$ generated by the PBT strategies, we execute the ground-truth function $f(x_{i})$ to deterministically obtain the expected output. This process yields a collection of $(x_{i},f(x_{i}))$ pairs, which form the rigorous test cases for the benchmark instance. Crucially, we incorporate a quality control step: only generated test suites that achieve high branch coverage (e.g., 100%) when executed on the ground-truth function are retained. We save a fixed number of high-coverage passing test cases (e.g., 500 pairs) for each problem. This generation process produces language-agnostic test cases for SC problems consisting of only basic types and library-aware test cases for WSC problems.
Testcase Runner Generation
For each benchmark instance and target evaluation program language, a Testcase Runner script is generated. The Runner’s primary role is to load PBT input-output pairs, execute the LLM-generated code, and perform differential testing [34] by comparing outputs against the ground truth. For SC tasks, the Runner is language-specific, designed to execute the language-agnostic test cases in the target language. For WSC tasks, the Runner is library-specific, incorporating utility functions to handle library-specific data types (e.g., numpy.allclose for arrays). Across all instances, a predefined recursive deep comparison function ensures accurate output validation. A dry run is performed during benchmark construction to verify the correctness of the generated Runner.
Instruction Generation
To provide clear and accurate guidance for LLM, we construct a comprehensive instruction for each benchmark instance. Inspired by the format and docstring generation process used in R2E [35], our instruction template includes a general directive to the LLM, followed by the function signature and its corresponding docstring embedded within a code snippet block. We use an LLM (distinct from the evaluated models) to refine the original function docstring or generate one if missing. To enhance robustness and mitigate potential biases from the original or LLM-generated text, inspired by prior work [23, 36], we employ back-translation using NL-Augmenter [37] to perturb the docstrings. Crucially, docstrings are tailored to task type: language-agnostic for SC tasks to support cross-language evaluation, and lib-aware for WSC tasks to precisely specify interactions with allowed libraries using relevant types.
3 The Code2Bench-2505 Benchmark
Code2Bench-2505 is the first benchmark instance constructed using Code2Bench, focusing on Python as the primary language for problem extraction while supporting cross-language evaluation for SC tasks. Dynamically sourced from high-quality open-source repositories, it ensures relevance and mitigates training data contamination risks. We sampled 880 Python projects on GitHub with at least 500 stars, targeting mature and well-maintained codebases. To ensure data lies beyond the training scope of most existing models (knowledge cutoff: August 2024) [38, 39], we extracted function candidates from commits between August 2024 and May 2025. Through Code2Bench’s multi-stage filtering pipeline, an initial pool of 1,442 candidates was refined into a final benchmark dataset of 1,163 instances. More details and examples can be found in Appendix H. The Code2Bench-2505 exhibits the following characteristics:
  • Function Category and Complexity Distribution: The dataset includes 965 Self-Contained (SC) tasks and 198 Weakly Self-Contained (WSC) tasks, reflecting varying levels of external dependency. SC tasks exhibit higher structural complexity than WSC tasks, with average cyclomatic complexity of 5.21 vs. 2.23 and average Lines of Code (LoC) of 21.25 vs. 17.25. Tasks are classified into difficulty levels (Easy, Medium, Hard) based on a capable LLM’s judgment, considering factors beyond structural metrics to reflect perceived complexity. For SC tasks, the distribution is Easy: 69, Medium: 86, Hard: 38; for WSC tasks, it is Easy: 75, Medium: 90, Hard: 33. This diverse difficulty distribution ensures comprehensive evaluation across real-world coding challenges.
  • Test Suite Quality: For each instance, we generated an average of 480 test cases for SC, 456 for WSC. These test cases achieve an average branch coverage of 100% on the ground-truth implementations, demonstrating the comprehensiveness and rigor of our test suite.
  • Problem Diversity and Dependencies: SC problems span a wide range of programming domains, including Data Processing, Mathematics, File I/O, Time Handling, Image Utilities, Logic Control, ensuring comprehensive coverage of real-world software tasks. WSC problems extend this scope by incorporating dependencies from a controlled set of 37 external libraries, such as numpy, pandas, re, and collections, reflecting realistic usage patterns in practical development scenarios.

### v2

# Code2Bench: Scaling Source and Rigor for Dynamic Benchmark Construction

@@ Abstract
The evaluation of code-generating Large Language Models (LLMs) is fundamentally constrained by two intertwined challenges: a reliance on static, easily contaminated problem sources and the use of superficial, low-rigor testing. This paper introduces a new benchmark construction philosophy, Dual Scaling, designed to systematically address both limitations. Our approach involves continuously scaling the source of problems from dynamic, real-world code repositories and systematically scaling the rigor of tests via automated, high-coverage Property-Based Testing (PBT). We instantiate this philosophy in Code2Bench, an end-to-end framework that leverages Scope Graph analysis for principled dependency classification and a 100% branch coverage quality gate to ensure test suite integrity. Using this framework, we construct Code2Bench-2509, a new benchmark suite with native instances in both Python and Java. Our extensive evaluation of 10 state-of-the-art LLMs on Code2Bench-2509, powered by a novel "diagnostic fingerprint" visualization, yields three key insights: (1) models exhibit a fundamental performance gap, excelling at API application (Weakly Self-Contained tasks) but struggling with algorithmic synthesis (Self-Contained tasks); (2) a model’s performance is profoundly shaped by the target language’s ecosystem, a nuance we are the first to systematically quantify; and (3) our rigorous, scaled testing is critical in uncovering an "illusion of correctness" prevalent in simpler benchmarks. Our work presents a robust, scalable, and diagnostic paradigm for the next generation of LLM evaluation in software engineering. The code, data, and results are available at https://code2bench.github.io/.
@@ 1 Introduction
As Large Language Models (LLMs) are increasingly integrated into software development workflows C. E. Jimenez, J. Yang, A. Wettig, S. Yao, K. Pei, O. Press, and K. Narasimhan (2023); 3; 1, the need for accurate and realistic evaluation of their coding capabilities has become paramount. However, the current landscape of code benchmarks is fundamentally constrained by two intertwined challenges: a reliance on static, easily contaminated problem sources and the use of superficial, low-rigor testing.
First, ❶ the static nature of canonical benchmarks like HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) leads to an inevitable obsolescence; their problems, having existed for years, are likely part of LLM training corpora, turning evaluation into an exercise in memorization rather than true generalization Carlini et al. (2021); Sainz et al. (2023). While dynamic “live” benchmarks Jain et al. (2024a); Li et al. (2024b) have emerged, they often source problems from competitive programming, which may not reflect the complexity of real-world software engineering. Second, ❷ the superficial testing common to most benchmarks, often relying on a handful of example-based tests, creates an illusion of correctness. As highlighted by EvalPlus Liu et al. (2023a), this insufficient test rigor fundamentally limits their ability to uncover the subtle, edge-case failures that define the gap between functional code and production-ready software. As summarized in Table 1, existing methods fall short across key dimensions, highlighting the significant limitations remaining and the necessity to design new benchmarks. To break this cycle of obsolescence and superficiality, we argue that a paradigm shift is needed. We propose a new benchmark construction philosophy centered on two core principles: (1) Scaling the Source, by dynamically and continuously ingesting a diverse array of problems from the ever-evolving landscape of real-world code repositories; and (2) Scaling the Rigor, by systematically generating comprehensive test suites with deep, verifiable coverage through Property-Based Testing (PBT) Claessen and Hughes (2000).
We instantiate this philosophy in CODE2BENCH, a novel, end-to-end framework that automates this dual-scaling process. ❶ To Scale the Source, CODE2BENCH first addresses the challenge of classifying diverse, real-world code. Our analysis of the existing benchmark landscape reveals an implicit bifurcation in evaluation focus, which we formalize into two primary task categories: (1) Self-Contained (SC) tasks, which require pure, dependency-free logic, reflecting the focus of benchmarks like HumanEval Chen et al. (2021) on core algorithmic reasoning; and (2) Weakly Self-Contained (WSC) tasks, which require the correct application of common libraries, capturing the focus of benchmarks like BigCodeBench Zhuo et al. (2024) on practical API application. This principled classification, enabled by our Scope Graph-based analysis, allows us to systematically generate tasks that target these distinct developer skills. ❷ To Scale the Rigor, CODE2BENCH then employs a powerful Property-Based Testing (PBT) engine and a stringent 100% branch coverage quality gate. This ensures that every problem in our benchmark is not only realistic but also a fully-explorable logical challenge, backed by a test suite capable of deep, diagnostic validation.
Using this framework, we construct CODE2BENCH-2509, a new, multi-faceted benchmark suite with native instances in Python and Java, curated from recent, real-world repositories. Our extensive evaluation of 10 state-of-the-art LLMs on this suite demonstrates the power of our approach. The synergy of a scaled source and scaled testing enables an unprecedentedly fine-grained diagnostic, revealing: (1) a performance gap between models’ ability in algorithmic synthesis (SC) and API application (WSC); (2) the profound impact of language paradigms on model failure modes, a nuance we are the first to systematically quantify; and (3) the critical role of rigorous testing in uncovering the “illusion of correctness” prevalent in simpler benchmarks.
We summarize our contributions as follows:
  • We propose CODE2BENCH, a novel framework that introduces and operationalizes the Dual Scaling philosophy for benchmark construction, systematically scaling the source of problems with dynamic acquisition and the rigor of tests with a 100% coverage PBT quality gate.
  • We construct and release CODE2BENCH-2509, a high-quality, contamination-resistant benchmark with native tasks in Python and Java, demonstrating significantly higher complexity and test rigor than prior work (Table 1).
  • We provide a deep, diagnostic analysis of state-of-the-art LLMs, introducing novel visualizations and uncovering key insights into their strengths and weaknesses in real-world coding scenarios.
@@
Table 1: Comparison of Code2Bench-2509 with existing code generation benchmarks.
@@
HumanEval Chen et al. (2021) Manual
MBPP Austin et al. (2021) Manual
EvalPlus Liu et al. (2023a) Manual
LiveCodeBench Jain et al. (2024a) Contests
RepoBench Liu et al. (2023b) Project Codebases
HumanEval-X Zheng et al. (2023b) Manual
BigCodeBench Zhuo et al. (2024) Synthetic
DevEval Li et al. (2024c) Project Codebases
EvoCodeBench Li et al. (2024b) Project Codebases
Code2Bench-2509 (Ours) Project Codebases
2 The Code2Bench Framework: A Dual Scaling Approach
@@
In this section, we detail the architecture and technical components of the Code2Bench framework. Our methodology is built upon the core philosophy of Dual Scaling: continuously scaling the source of benchmark problems to ensure realism and novelty, and systematically scaling the rigor of our tests to enable deep, diagnostic evaluation. We organize our discussion around these two principles.
2.1 Scaling the Source: Dynamic Acquisition from Real-World Code
Temporal Filtering for Contamination Resistance. The primary threat to the validity of LLM evaluation is data contamination, where benchmarks become obsolete as their contents are absorbed into training corpora. To combat this inevitable obsolescence, our framework’s first principle is to dynamically source problems that are provably unseen. We achieve this through Temporal Filtering, a deterministic strategy grounded in the version control timestamps of real-world code. Our method leverages a simple axiom: a model cannot have been trained on code that did not exist prior to its knowledge cutoff date. For each model under evaluation, we run our acquisition pipeline to extract functions exclusively from GitHub commits created after its official knowledge cutoff.
Scope Graph-based Analysis for Dependency Classification.
To impose a meaningful structure on our diverse pool of functions, we automate their classification based on dependencies. We employ a Scope Graph-based analysis Néron et al. (2015)—a formal, language-agnostic method that precisely identifies all external dependencies, a task where simpler methods like AST traversal often fail.
Our classification algorithm is a deterministic, two-step process: (1) Dependency Identification, where we use the Scope Graph to compute the set of all unresolved references ($\mathcal{D}$) for each function. (2) Rule-based Classification, where we apply rules based on a predefined allowed libraries, $\mathcal{L}_{\text{allowed}}$:
  • If $\mathcal{D}=\emptyset$, it is classified as Self-Contained (SC) (pure algorithmic reasoning).
  • If $\mathcal{D}$’s dependencies resolve entirely within $\mathcal{L}_{\text{allowed}}$, it is Weakly Self-Contained (WSC) (API application).
  • Otherwise, it is discarded (e.g., Project-Dependent).
This principled, automated process is a cornerstone of our framework, enabling the targeted evaluation of distinct model capabilities.
Program Analysis for Testability and Complexity.
Following dependency classification, we apply a final layer of automated program analysis to ensure candidates are both testable and non-trivial. First, to guarantee testability, we use Control-Flow Graph (CFG) analysis to discard functions lacking a verifiable, input-dependent output (e.g., no return statement). Second, to ensure a meaningful complexity, we filter functions based on their Cyclomatic Complexity McCabe (1976), targeting a range (e.g., $[2,10]$) that balances challenge with solvability. This dual-filtering stage is crucial for curating a high-quality benchmark of tasks that are both evaluable and diagnostically valuable.
LLM-based Semantic Filtering.
While prior analyses ensure structural soundness, they cannot distinguish a meaningful task from a trivial one. To assess for semantic relevance and conceptual challenge, we therefore employ an LLM-as-a-Judge Zheng et al. (2023a); Li et al. (2024a). This final filtering step, designed for high reliability via deterministic decoding and a structured classification prompt, ensures our benchmark contains problems of genuine substance. A rigorous validation, detailed in Appendix D, confirms its near-perfect agreement with human experts (Cohen’s $\kappa=0.95$).
2.2 Scaling the Rigor: Automated Synthesis via Property-Based Testing(PBT)
Property-Based Testing (PBT) for Comprehensive Input Generation.
Traditional example-based tests verify a function against a small, fixed set of known inputs. In contrast, Property-Based Testing (PBT) Claessen and Hughes (2000) explores a much larger input space by generating hundreds or thousands of random, yet structured, inputs and asserting that a general property of the code holds true for all of them. In our framework, the core property we test is functional equivalence with the ground-truth implementation. For any valid input $\mathbf{x}$ generated by our PBT engine, the output of an LLM-generated function $f_{\text{LLM}}(\mathbf{x})$ must match the output of the original, real-world ground-truth function $f_{\text{gt}}(\mathbf{x})$. The ground-truth function thus serves as a perfect test oracle.
The process of input synthesis is driven by automated strategy generation. For each function candidate, our framework analyzes its signature, including parameter types and type hints, to compose a set of PBT strategies. These strategies are not simple random generators, but intelligent explorers of the input domain, designed to produce a rich distribution of values—including typical inputs, boundary cases (e.g., empty lists, zeros, min/max values), and complex nested structures (e.g., variable-shaped lists of dictionaries). This automated process yields a comprehensive suite of hundreds of input-output pairs $(\mathbf{x}_{i},f_{\text{gt}}(\mathbf{x}_{i}))$ for each function, forming the foundation for the rigorous validation described next. The specific PBT libraries used for each language (e.g., Hypothesis for Python, jqwik for Java) are detailed in Appendix E.
The “Great Filter”: A 100% Coverage Quality Gate.
While Property-Based Testing generates a high volume of diverse inputs, quantity alone does not guarantee rigor. A test suite, however large, is only effective if it thoroughly exercises the internal logic of the function under test. To enforce this level of rigor systematically, we introduce the final and most stringent stage of our pipeline: the “Great Filter”, a quality gate that mandates 100% branch coverage.
The mechanism is as follows: after a PBT suite is synthesized for a ground-truth function $f_{\text{gt}}$, we execute the entire suite against $f_{\text{gt}}$ itself and measure the resulting branch coverage using standard language-specific tools (e.g., coverage.py). A function candidate and its corresponding test suite are only accepted into the final benchmark if and only if this execution achieves 100% branch coverage. This seemingly simple requirement has a profound impact on the final benchmark’s quality and character. It acts as a powerful, dual-purpose filter:
  • It filters out inadequate tests. If a PBT suite fails to achieve full coverage, it indicates that the input generation strategy was not sophisticated enough to explore all logical paths of the function. Such a test suite would be incapable of providing a truly rigorous evaluation, and is discarded.
  • It filters out untestable functions. More importantly, if even a well-designed PBT strategy cannot trigger all branches, it often signals that the function itself is "untestable" in isolation. This typically occurs in functions with defensive code for unreachable states, complex error handling coupled to external systems, or other logic that cannot be exercised through its public API. These functions, while present in real-world code, are unsuitable for a standalone, functional correctness benchmark.
The “Great Filter” is the primary reason for the significant reduction in candidates observed in our data funnel (as shown in Figure 1). It is a deliberate trade-off, prioritizing uncompromising rigor over sheer volume. The result is a smaller, but significantly more potent, benchmark where every single problem is guaranteed to be a non-trivial, fully-explorable logical puzzle, backed by a test suite capable of validating every branch of its solution.
Instruction Generation for Task Specification.
To ensure a fair and effective evaluation, each task is accompanied by a clear, unambiguous instruction. Our framework automates the generation of these instructions by refining a function’s original source docstring and signature using a powerful LLM (GPT-4o) with deterministic decoding. To mitigate potential biases, we also employ a back-translation perturbation technique Zhuo et al. (2024); Wang et al. (2022); Dhole et al. (2021).
Crucially, the instruction style is systematically adapted to the task’s dependency classification, ensuring the model is provided with the precise context needed for the specific challenge:
  • For SC Tasks(e.g., SC-Python, SC-Java): Instructions use language-native conventions—Python docstrings with types like list and dict for SC-Python, and Javadoc with Java types like List<String> for SC-Java. Though library-free, these tasks assess a model’s proficiency with each language’s core built-in features and data structures.
  • For WSC Tasks: The instruction is made library-aware and language-native. It explicitly names the required external libraries (e.g., NumPy) and uses the precise, idiomatic types of the target language’s ecosystem (e.g., numpy.ndarray). This targets the evaluation on a model’s practical ability to correctly apply common APIs.
Benchmark Instantiation and Runner Generation. The final step of our framework packages each curated problem into an executable benchmark instance. This instance comprises two key components: a test suite and a test runner. The test suite, containing hundreds of input-output pairs, is generated using a native Property-Based Testing (PBT) engine (e.g., Hypothesis, jqwik) to ensure high coverage and rigor. To conduct the evaluation, a corresponding language-native Test Runner is automatically generated. The runner is responsible for deserializing the test suite, executing the LLM-generated code, and performing a rigorous deep comparison against the ground-truth outputs. To guarantee the runner’s correctness, we perform a dry run where the LLM’s function is replaced by the ground-truth function, ensuring the entire test harness passes flawlessly before evaluation. This end-to-end native approach, validated by a dry run, ensures our evaluation is both stringent and reliable. Further details are in Appendix F.
3 The Code2Bench-2509 Benchmark Suite
Table 2: Quantitative characteristics of the CODE2BENCH-2509, compared to prior benchmarks.
Metric / Dimension SC-Python WSC-Python SC-Java HumanEval MBPP
I. Scale & Complexity
   # Tasks 217 194 249 164 974
   Avg. Lines of Code (LoC) 20.6 18.3 14.1 7.3 6.5
   Avg. Cyclomatic Complexity (CC) 5.3 2.6 3.6 2.8 2.3
   Difficulty (E:M:H Ratio) 0.30:0.40:0.30 0.28:0.41:0.31 0.27:0.43:0.30 - -
II. Testing Rigor
   Avg. Test Cases per Task ~500 ~500 ~500 ~7.8 ~3.0
   Test Coverage Guarantee 100% Branch 100% Branch 100% Branch Variable Variable
III. Diversity & Extensibility
   Source Type Real-World Real-World Real-World Hand-Crafted Crowd-Sourced
   Dependency Scope Self-Contained >30 Libraries Self-Contained Self-Contained Self-Contained
   Language Extensibility (Python) (Python) Java Native (Python) (Python)
Figure 2: The CODE2BENCH multi-dimensional evaluation landscape.
Code2Bench-2509 is a new benchmark suite, automatically curated from May to September 2025, designed to overcome the limitations of prior benchmarks by systematically expanding evaluation along three key dimensions: Testing Rigor, Dependency Level, and Framework Extensibility. Figure 2 visually situates our benchmark within this landscape, showcasing its significant leap forward compared to predecessors like HumanEval and BigCodeBench. We targeted actively maintained, open-source repositories on GitHub. To minimize noise and bias, we enforced the following criteria: (1) Community Validation: Repositories must have $\geq 500$ stars; (2) Active Maintenance: Commits within the last 3 months; and (3) Domain Diversity: We employed stratified sampling across 10 diverse domains (e.g., Web, ML, System) to prevent over-representation of any single field. We actively filtered out homework assignments and tutorials. A detailed disclosure of the selection criteria, sampling strategy, and the full repository list is provided in Appendix H. The quantitative evidence for this advancement is detailed in Table 2. Our instances demonstrate significantly higher structural complexity (e.g., an average Cyclomatic Complexity of 5.3 for SC-Python vs. 2.8 for HumanEval) and an order-of-magnitude increase in testing rigor, featuring ~500 test cases per task with a guaranteed 100% branch coverage. Beyond these metrics, the suite’s high quality is rooted in the rich diversity of its tasks, sourced from 220 Python and 189 Java recent, real-world repositories. This ensures wide topical coverage, while the successful instantiation of a native Java suite provides concrete proof of our framework’s extensibility. This combination of complexity, rigor, diversity, and extensibility provides a more challenging and realistic platform for assessing the true capabilities of modern LLMs. More details can be found in Appendix H.